The period of a pendulum on two planets are in the ratio 3:4the acceleration due to gravity on them are in ratio
 
 

Grade:11

2 Answers

Arun
4 years ago
 
Dear student 
 
As we know T = 2 pi sqrt (L/g)
 
hence
 
T1 /T2 = sqrt (g2/g1)
 
¾ = sqrt (g2/g1)
 
hence g1 / g2 = 16/9
Vikas TU
4 years ago
Dear student 
Timeperiod of pendulum = T = 2pie√(L/g)
T^2 is directly proportional to 1/g 
g is directly proportional to 1/T^2 
g1/g2 = (T2/T1)^2 
g1/g2 = 16/9 
Hope this helps